OpenStudy (anonymous):

A gardener wants to create a triangular garden in the shape of a right triangle with shortest side length x – 2y ft. and the middle length side x + 3y ft. What is an algebraic expression for the area of the garden? Be sure to multiply this out and express in simplest correct mathematical form, including units.

OpenStudy (anonymous):

a = 1/2(base)*(height), so: 1/2(x-2y ft)*(x+3y ft) = (x^2 -6y^2 + xy)/2 ft^2

OpenStudy (anonymous):

is that the answer??

OpenStudy (shadowfiend):

Yep!

OpenStudy (anonymous):

im not understanding how you got that

OpenStudy (shadowfiend):

It was three parts. First off, if you draw a right triangle, you'll see that depending on which direction you draw it in, either side can be the `base' or the `height'. Second, you have the equation, \(\frac{1}{2}bh\), which is one half of the base times the height. That's the equation for the area of a triangle. Lastly, you plug the two sides they give you in for the base and height, so you get: \[\frac{1}{2}(x - 2y) (x + 3y)\] And then you multiply all of that out, which is what gives you the result nothingmaster came up with.

OpenStudy (anonymous):

ok....i get the 1/2*bh but im not sure how that answer comes out.....I SUCK ATMATH!!! lol math makes me wanna cry